# Minimum Cost Loop Nests for Contraction of a Sparse Tensor with a Tensor Network

### v1

@@ 8. Conclusion and Future Work
Favorable performance of SpTTN-Cyclops in comparison to other general tensor contraction libraries, as well as comparisons to specialized codes, demonstrate that implementation of high-performance SpTTN kernels of interest to tensor decomposition and completion can be effectively automated. As opposed to prior frameworks for sparse tensor contractions, by restricting consideration to a single sparsity pattern and dense buffers, we are able to enumerate and efficiently find the minimum cost SpTTN loop nest. At the same time, the resulting implementations are practical, as they may be accelerated by standard BLAS libraries, and match the structure of existing optimized codes specialized to particular SpTTN contractions. Our framework and evaluation of SpTTN kernels can be extended in several ways. For example, the search space can be extended to include partially-fused loop nests. This would be particularly useful for cost metrics like number of BLAS kernels or the degree of parallelism when fused terms are independent. In addition, more sophisticated cost models can be developed, and the evaluation of approximate auto-tuning techniques based on these could be explored.
@@ Acknowledgements.
This research has been supported by funding from the United States National Science Foundation (NSF) via grants #1942995 and #1931258, as well as by the Department of Energy (DOE) Advanced Scientific Computing Research program via award DE-SC0023483. Raghavendra Kanakagiri has been supported by NSF grant #1931258 and the University of Illinois Urbana-Champaign Computer Science Future Faculty Fellows program. This work used Stampede2 at TACC through allocation
CCR180006 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.

### v2

@@ 8. Conclusion and Future Work
Favorable performance of SpTTN-Cyclops in comparison to other general tensor contraction libraries, as well as comparisons to specialized codes, demonstrate that implementation of high-performance SpTTN kernels of interest to tensor decomposition and completion can be effectively automated. As opposed to prior frameworks for sparse tensor contractions, by restricting consideration to a single sparsity pattern and dense buffers, we are able to enumerate and efficiently find the minimum cost SpTTN loop nest. At the same time, the resulting implementations are practical, as they may be accelerated by standard BLAS libraries, and match the structure of existing optimized codes specialized to particular SpTTN contractions. Our framework and evaluation of SpTTN kernels can be extended in several ways. For example, the search space can be extended to include partially-fused loop nests, which may offer additional parallelism.
@@ Acknowledgements.
This research has been supported by funding from the United States National Science Foundation (NSF) via grants #1942995 and #1931258, as well as by the Department of Energy (DOE) Advanced Scientific Computing Research program via award DE-SC0023483. Raghavendra Kanakagiri has been supported by NSF grant #1931258 and the University of Illinois Urbana-Champaign Computer Science Future Faculty Fellows program. This work used Stampede2 at TACC through allocation CCR180006 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.